# Privacy risk in GeoData: A survey

### v1

@@ 2.2.4. Differencing attack
A differencing attack is possible with variables for which there are multiple different plausible coding schemes, where the categories in those coding schemes are not nested but instead overlap. For example, given a table with information on 20–25 year olds and a table with information on 20–24 year olds, then the difference between the two tables will reveal information about 25 year olds only. If the difference between cells contains a small number of individuals then a disclosure is more likely.
Differencing attacks occur when data released at different levels of geography can be combined to reconstruct data at a smaller level or identify the location of an observation. This situation may occur for tables or maps with different geographical codings potentially allowing more information to be revealed about individuals in the overlaps than intended from each individual table. Although it could happen with any variable the issue most commonly comes up with geography. The result of this differencing is that whilst a table may be considered safe in isolation, this may not be the case for multiple tables when overlain with one another.
@@ 3. A Taxonomy of Geomasking
In this section we describe a taxonomy for geomasking techniques. Our aim in developing this taxonomy is to provide a clearer picture of current approaches to geomasking, and to identify gaps in these techniques which will help us to identify directions for future research. We describe 11 dimensions of geomasking which we categorise into three main topics, as is illustrated in Fig. 2. Combined, these 11 dimensions provide a comprehensive characterisation of geomasking techniques. In the following sub-sections we discuss each dimension in detail, and we provide an overview of the major methodologies or techniques applied in these dimensions.
Figure 2. Classification of the surveyed research on privacy preserving geomasking techniques.
@@ 3.1. Privacy techniques for Geomasking
Geomasking is a class of methods for changing the geographic location of an individual in an unpredictable way to protect confidentiality, while trying to preserve the relationship between geocoded locations and occurrence events. In this section we analyse different privacy techniques proposed for geomasking.
Following earlier work on geomasking (Stinchcomb, 2004), we group geomasking techniques into five major categories: (1) random perturbation, (2) masking with pre-defined locations, (3) affine transformation, (4) aggregation and clustering, and (5) Synthetic data generation . In the next subsections we discuss different techniques that are proposed under each of these categories in more details.
@@ 3.1.1. Random Perturbation
Random perturbation geomasking moves each individual points from their original location to a random distance in a random direction (Allshouse et al., 2010). When the generated masked points may be close to the original points and for a small proportion of the geomasked points, the adversary can re-identify individuals by reversing geocode. Next, we explain several techniques that can categorised under random perturbation.
@@ 3.1.1.1 Density-adaptive displacement
In this category, the displacement distance is inversely proportional to the underlying population density. In the geomasking literature the population densities can be considered either as uniform or heterogeneous.
@@
Uniform population density distribution: The methods under this category assume the population density is distributed homogeneously. Figure 3 illustrates the different method under this category.
@@
    Donut Masking (random perturbation within an annulus): The donut method extends current methods of random displacement by ensuring that an address is not randomly assigned on or too near its original location. In donut method geomasking, each geocoded address is relocated in a random direction by at least a minimum distance, but less than a maximum distance. In addition, each point is moved a distance inversely proportional to the underlying population density, which provides privacy protection while minimising the introduced spatial error. For example, persons in high-density urban areas do not need to be moved as far as persons in low-density rural areas to achieve the same magnitude of anonymisation. The donut method is an adaptive geomask because the dimensions of the mask around each point vary to meet specified anonymity constraints based on the underlying population density. However, in cases where population density is not evenly distributed, there may be an under-protection of individuals (Allshouse et al., 2010).
@@
    Spatially adaptive random perturbation (SARP): This technique considers the spatial variation of the population at risk when displacing points by random distances and in random directions (Lu et al., 2012). It reduces unnecessary noise in high-dense population zones by using relatively large and small population zones for low and high population densities, respectively. The SARP geomasking technique defines the perturbation zone based on the actual distribution of residential addresses (risk location) instead of people (risk population). It employs ”donut-shaped” spatially varied perturbation zones rather than ”pancake-shaped” which allows a point to be moved to a nearby location within its immediate surroundings. Utilising actual street addresses the SARP geomasking technique could offer better privacy protection compared to geomasking based on population size. While the SARP techniques do not significantly alter clustering patterns on a global scale, the geomasked data tends to exhibit greater clustering compared to the real distribution.
@@
    Weighted random perturbation (WRP): Weighted random perturbation is included because of its potential to better preserve spatial distributions by displacing points shorter distances in more populated areas (Kwan et al., 2004b). The weighted random mask aimlessly rotates every data point on a circle applying the variable radius $r$, where $r$ depends on the underlying population density. This approach enhances privacy protection by mitigating the risk of disclosing spatial information. Additionally, it imposes a maximum displacement limit based on the distance to each point’s $k$-th nearest neighbour.
@@ 3.1.1.2 Gaussian deviation for displacement magnitude
    Density–based Gaussian spatial skew: In this method the location points are adjusted by applying a random offset derived from a Gaussian distribution, where the standard deviations of the distribution are inversely related to the population density in the local area (Cassa et al., 2006). Utilising local demographic data allows the proposed anonymisation system to shift individuals in heavily populated regions by shorter distances compared to those in sparse areas. Consequently, addresses can undergo minimal distortion while upholding a designated level of k-anonymity.
@@ 3.1.1.3 Multiple risk factor adaptive displacement
This category incorporates various factors in calculating displacement distances, including but not limited to population density. Triangular Displacement is a method that can be considered under this category. Triangular Displacement method considers population density for displacement distances, incorporates multiple factors, such as individual re-identification risk, data sensitivity levels, research types, the presence of quasi-identifiers, and potential exposure to external sources, to calculate displacement distances (Murad et al., 2014). However, this technique has limitations. Triangular Displacement method may be vulnerable to edge effects and could result in masked locations falling into impractical or unrealistic areas, thereby impacting data confidentiality. Despite these limitations, this method represents a step forward in geomasking techniques, offering a more nuanced and context-aware approach to safeguarding spatial data in research and analysis.
@@ 3.1.2. Masking with pre-defined locations
The proposed methods under this category move a given point to a new location, which is randomly chosen from all the possible locations of a similar type that exist within a specific region (i.e., a circle or an annulus) surrounding the original location.
3.1.2.1 Location Swapping
@@
Location swapping method swaps a given location/point to a new location based on the similar geographic characteristics of a point within a specified neighbourhood (Zhang et al., 2017). The method first predefined a radius around the location to be swapped. Then the method randomly selects a location among all the possible locations that falls within the defined radius. This method is similar to the random perturbation within a circle due to couple of reasons: (1) Within the predefined radius all the locations are considered with equal probability making the swapping to any location possible within the defined radius and (2) the radius can be defined based on the underlying population density distributions. However, the location swapping method considered only the locations that have predefined locations (such as residential addresses) as possible candidates for swapping and the radii can be varied based on the local population densities. Figure 4 (right plot) illustrates an example of location swapping.
3.1.2.2 Location Swapping within Donut
@@
This method represents another variation of location swapping, employing the concept of donut masking to set a minimum distance, preventing points from being swapped with any locations within this specified range (Zhang et al., 2017). The determination of this minimum distance can be based on local population densities. The authors employed an internal buffer size half that of the external buffer.
3.1.2.3 Street Masking
@@
Similar to Location swapping, in the street masking method the original location is replaced with coordinates of a point sampled from a set of points within a predefined radius (Swanlund et al., 2020b). However, for sampling this method uses road networks from a given map instead of using residential addresses. However, when selecting points, points that are not close to a drivable road (such as in rural areas) are additionally moved, as all points are first moved to the nearest road that is applicable. Similar to location swapping, street masking can utilise the population densities when determining the radius assuming that areas with a high population density also have high road density and vise versa. However, for high road density areas this method could consume larger runtime due to the number of points to be masked.
@@ 3.1.3. Affine-Transformation
The transformations techniques, which is also known as isomasks techniques, apply basic geometric transformations to move spatial point data (Ajayakumar et al., 2019). Here, it can be differentiated between rotate, scale and translate, stochastic mechanism, and concentration of isomasks. Affine transformations change the original location by either twisting every point by a set angle (referred to as rotation), or moving every point by a set increment (referred to as translation) or by moving every point by scaling constant (referred to as scale). Next, we describe these methods in more details and Fig 5 illustrates some of these methods.
3.1.3.1 Translation
This method adds a random constant displacement value to both x and y coordinates while ensuring the distance between coordinates remains the same (Wang et al., 2022). All points will be shifted to a fixed distance and direction.
3.1.3.2 Flipping
This method manipulates the orientation of axes, both on a local and global scale. There are several flipping techniques, such as global horizontal flipping, which flips the map’s horizontal axis, global vertical flipping, which flips the map’s vertical axis, global axes flipping, which flips both horizontal and vertical axes, and local random flipping, which randomly alters the orientation of horizontal, vertical, or both axes within grid cells to hide location points (Gupta and Rao, 2020; Leitner and Curtis, 2004, 2006).
3.1.3.3 Rotating
This method rotates each point by a fixed angle ($0\leq\theta\leq 360$), centering the rotation around either the origin of the coordinate system or an arbitrarily chosen point (Wang et al., 2022).
3.1.3.4 Scaling (Resizing)
This method multiplies a constant value with both x and y coordinates, which remains consistent for both axes (Wang et al., 2022). Consequently, this process alters the distances between the points and all points will be expanded by a scaling factor.
3.1.3.5 Horizontal shear
This method employs a linear transformation to horizontally shear the location points (Armstrong et al., 1999; Broen et al., 2021). Each point is shifted along its x-axis, oriented at a 45° angle relative to its initial position concerning the center of the point distribution.
@@ 3.1.4. Spatial aggregation & Clustering
Spatial aggregation aggregates individual location points within specific geographic units or areas, such as cities, zip codes, or administrative areas, to ensure anonymity and prevent the disclosure of exact locations. This technique operates by decreasing the spatial resolution of the data, thereby modifying the original data patterns. The goal is to generate generalised records that consolidate information within defined territories, protecting the identities of specific locations. However, this process may compromise the utility of published records as the detailed patterns found in high-resolution data might not be as discernible in the aggregated information.
Clustering aims to reveal inherent patterns within spatial data by grouping similar data points based on geographical proximity or shared characteristics. Clustering maintains the overall characteristics of an area without pinpointing precise locations, preserving the general patterns while ensuring individual data privacy. However, it’s essential to note that the scalability and computational complexity of clustering algorithms can present challenges, especially when handling large and diverse spatial datasets. Figure 6 illustrates several technique under this category.
@@ 3.1.4.1 Grid-based aggregations
The techniques in this category involves the realignment of individual data points onto a grid system (Leitner and Curtis, 2004). It typically involves relocating each original data point to specific positions within designated grid cells. These relocations may take place at either the centroid or vertices of the grid cells, each being aligned with a defined side length. The efficacy of methods within this category is notably influenced by two crucial parameters, (1) the size of the aggregation units and (2) the boundaries utilised to define these units.
In grid based aggregations the grid cell size could be fixed or varied. Grid line masking and grid center masking are two popular methods that use fixed size grid cells. Grid line masking method relocates location points to the nearest edge of the grid cell in which they are contained (Gupta and Rao, 2020). Grid center masking method relocates the location points to the centroid of the cell in which they are presently situated (Leitner and Curtis, 2004).
@@
[…]
@@
    Military Grid Reference System (MGRS) protects sensitive point-based geospatial data with a focus on preventing the matching of points with their true underlying observations, such as street addresses. This method involves converting coordinates to the Military Grid Reference System (MGRS) and using digit switching for masking, allowing encryption at different spatial precision levels (Clarke, 2016). The selection of the digit switching combination minimises differences in aggregate descriptive spatial statistics between the masked and unmasked data, with emphasis on the nearest neighbour statistic. Despite the computing time required for processing, the method offers new possibilities for protecting sensitive geospatial data, overcoming disadvantages of geometric point masking methods, such as the lack of control over parameters, compromised spatial analysis of distribution, and susceptibility to violations of geoprivacy associated with the use of high-resolution imagery and accurate data from Global Navigation Satellite Systems (GNSS). The proposed digit-switching method, which operates effectively across multiple nested spatial scales, addresses these limitations by allowing the analyst to choose a distribution that closely mimics the spatial distribution of unmasked data, ensuring suitability for spatial analysis based on aggregate descriptive statistics.

### v2

@@ 2.2.4. Differencing attack
A differencing attack is possible with variables for which there are multiple different plausible coding schemes, where the categories in those coding schemes are not nested but instead overlap. For example, given a table with information on 20- to 25-year-olds and a table with information on 20- to 24-year-olds, the difference between the two tables will reveal information about 25-year-olds only. If the difference between cells contains a small number of individuals, then a disclosure is more likely.
Differencing attacks occur when data is released according to a geospatial hierarchy such that data in different granularity levels can be combined to reconstruct data at a finer granularity or identify the location of an observation. This situation may occur for maps with different geographical coding potentially allowing more information to be revealed about individuals in the overlaps than intended from each individual table. Although it could happen with tabular data, the issue most commonly comes up with geodata due to the publicly available geospatial information. The result of this differencing is that whilst a single map may be considered safe in isolation, this may not be the case for multiple maps when overlain with one another.
@@ 3. A Taxonomy of Geomasking
In this section, we outline our proposed taxonomy encompassing three primary dimensions of geomasking techniques: privacy techniques, evaluation measures, and practical aspects, as illustrated in Fig. 2. As previously discussed in Section 1, our aim in developing this taxonomy is to thoroughly analyse existing geomasking techniques, the measures to evaluate their performance, potential datasets for experimental evaluations, existing purpose-built prototypes of the techniques, the diverse domains where these techniques can be applied, and to identify gaps in existing geomasking techniques. Such analysis will assist data custodians in navigating existing methods and prototypes to identify the most suitable ones based on their specific requirements. In the following subsections, we discuss each dimension in detail.
Figure 2. Taxonomy of geomasking.
@@ 3.1. Privacy techniques for Geomasking
Geomasking is a class of methods for changing the geographic location of an individual in an unpredictable way to protect confidentiality, while trying to preserve the relationship between geocoded locations and occurrence events. In this section, we analyse different privacy techniques proposed for geomasking. Following earlier studies (Armstrong et al., 1999; Stinchcomb, 2004; Zandbergen, 2014; Seidl et al., 2018; Kekana, 2020; Broen et al., 2021; Wang et al., 2022; Redlich, 2022; Kamel Boulos et al., 2022; Iyer et al., 2023), we group geomasking techniques into five major categories based on how each technique is treating the data points to provide privacy: (1) Random perturbation, (2) Masking with pre-defined locations, (3) Affine transformation, (4) Aggregation and clustering, and (5) Synthetic data generation. Next, we explore different techniques that are proposed under each of these categories and discuss their potential and limitations. The existing literature for each category is further discussed in Section 4.
@@ 3.1.1. Random Perturbation
Random perturbation moves individual points from their original location to a random distance in a random direction (Allshouse et al., 2010). When the generated masked points may be close to the original points and for a small proportion of the geomasked points, the adversary can re-identify individuals by reversing geocode. Next, we explain several techniques that can be categorised under random perturbation.
@@ 3.1.1.1 Density-adaptive displacement
In this category, the displacement distance is inversely proportional to the underlying population density. In the geomasking literature, the population densities can be considered either as uniform or heterogeneous.
@@
Uniform population density distribution: The methods under this category assume the population density is distributed homogeneously. Figure 3 illustrates the different methods under this category.
@@
    Donut Masking (random perturbation within an annulus): The donut method extends current methods of random displacement by ensuring that an address is not randomly assigned on or too near its original location. In donut masking, each geocoded address is relocated in a random direction by at least a minimum distance, but less than a maximum distance. In population-based donut masking, each point is moved at a distance inversely proportional to the underlying population density. This means that points within regions of lower population density are displaced over larger distances compared to those situated in areas of higher population density. Essentially, the displacement magnitude is dynamically adjusted based on the concentration of the population, ensuring that privacy protection is maintained while minimising spatial error. Consequently, points in sparsely populated regions undergo more substantial repositioning compared to points in densely populated urban areas, thereby striking a balance between data privacy and spatial accuracy (Stinchcomb, 2004; Allshouse et al., 2010).
@@
    Spatially adaptive random perturbation (SARP): This technique considers the spatial variation of the population at risk when displacing points by random distances and in random directions (Lu et al., 2012). It reduces unnecessary noise in high-dense population zones by using relatively large and small population zones for low and high population densities, respectively. The SARP geomasking technique defines the perturbation zone based on the actual distribution of residential addresses (risk location) instead of people (risk population). It employs ”donut-shaped” spatially varied perturbation zones rather than ”pancake-shaped” which allows a point to be moved to a nearby location within its immediate surroundings. Utilising actual street addresses the SARP geomasking technique could offer better privacy protection compared to geomasking based on population size.
@@
    Weighted random perturbation (WRP): This technique has the potential to preserve spatial distributions by displacing points shorter distances in more populated areas (Kwan et al., 2004b). The weighted random mask aimlessly rotates every data point on a circle, applying the variable radius $r$, where $r$ depends on the underlying population density. This approach imposes a maximum displacement limit based on the distance to each point’s $k$-th nearest neighbour.
@@ 3.1.1.2 Gaussian deviation for displacement magnitude
    Density–based Gaussian spatial skew: In this method, the location points are adjusted by applying a random offset derived from a Gaussian distribution, where the standard deviations of the distribution are inversely related to the population density in the local area (Cassa et al., 2006). Utilising local demographic data allows the proposed anonymisation system to shift individuals in heavily populated regions by shorter distances compared to those in sparse areas. Consequently, addresses can undergo minimal distortion while upholding a designated level of k-anonymity.
@@ 3.1.1.3 Multiple risk factor adaptive displacement
This category incorporates various factors in calculating displacement distances, including but not limited to population density. Triangular Displacement is a method that can be considered under this category. The triangular displacement method considers population density for displacement distances and incorporates multiple factors such as individual re-identification risk, data sensitivity levels, research types, the presence of quasi-identifiers, and potential exposure to external sources to calculate displacement distances (Murad et al., 2014). However, this technique has limitations. It may be vulnerable to edge effects and could result in masked locations falling into impractical or unrealistic areas, thereby impacting data confidentiality. Despite these limitations, this method represents a step forward in geomasking techniques, offering a more nuanced and context-aware approach to safeguarding spatial data in research and analysis.
@@ 3.1.2. Masking with pre-defined locations
The proposed methods under this category move a given point to a new location, which is randomly chosen from all the possible locations of a similar type that exist within a specific region (i.e., a circle or an annulus) surrounding the original location. Location swapping methods offer advantages over random methods by providing a more realistic approach to selecting displaced locations, ensuring these locations fall within predefined areas of interest rather than arbitrary placements (i.e., bodies of water or uninhabited are). This enhances anonymity and preserves spatial patterns more akin to unmasked locations. However, if a malicious actor can ascertain the buffer distance used for relocation, there remains a risk of re-identifying the original locations (Zhang et al., 2017).
  1. (1)
@@
    Location Swapping: This method swaps a given location/point to a new location based on the similar geographic characteristics of a point within a specified neighbourhood (Zhang et al., 2017). The method first predefined a radius around the location to be swapped. Then, the method randomly selects a location among all the possible locations that fall within the defined radius. This method is similar to the random perturbation within a circle for a couple of reasons: 1) Within the predefined radius, all the locations are considered with equal probability, making the swapping to any location possible within the defined radius, and 2) the radius can be defined based on the underlying population density distributions. However, the location swapping method considered only the locations that have predefined locations (such as residential addresses) as possible candidates for swapping, and the radii can be varied based on the local population densities. Figure 4 (right plot) illustrates an example of location swapping.
  2. (2)
@@
    Location Swapping within Donut: This method represents another variation of location swapping, employing the concept of donut masking to set a minimum distance, preventing points from being swapped with any locations within this specified range (Zhang et al., 2017). The determination of this minimum distance can be based on local population densities. The authors employed an internal buffer size half that of the external buffer.
  3. (3)
@@
    Street Masking: Similar to Location swapping, in the street masking method, the original location is replaced with coordinates of a point sampled from a set of points within a predefined radius (Swanlund et al., 2020b). Unlike location swapping, which relies on residential addresses, this method utilises road networks from a given map for sampling. During point selection, any points not in proximity to a drivable road (such as in rural areas) are further adjusted by moving them to the nearest applicable road. Similar to location swapping, street masking can leverage population densities to determine the radius, assuming areas with high population density also exhibit high road density and vice versa. Nonetheless, in areas with high road density, this method may experience longer runtimes due to the larger number of points to be masked.
Street masking offers a convenient solution for geographic masking by leveraging road network data, which is widely available globally. By aligning data points with street networks, street masking preserves topological continuity between original and masked locations and minimises false attribution risks, ensuring realistic shifts without arbitrary displacements. However, its reliance on road networks may introduce biases towards urban and built environments, potentially skewing analytical outcomes. In contrast, location swapping offers a viable alternative by allowing for more controlled displacement of data points without the inherent biases associated with street-centric masking. Location swapping is preferable in scenarios where avoiding point displacement onto roads is essential. This method overcomes issues related to urban biases and road network dependencies by swapping points with verified neighbors or alternative locations. Street masking lacks intuitive control over the search depth parameter, leading to varying masking distances influenced by local road configurations and density. This lack of control complicates the process of determining optimal settings, posing a challenge for users seeking consistent results across different datasets and study areas.
@@ 3.1.3. Affine-Transformation
The transformation techniques, which are also known as isomasks techniques, apply basic geometric transformations to move spatial point data (Ajayakumar et al., 2019). Here, it can be differentiated between rotate, scale and translate, stochastic mechanism, and concentration of isomasks. Affine transformations change the original location by either twisting every point by a set angle (referred to as rotation), moving every point by a set increment (referred to as translation) or moving every point by scaling constant (referred to as scale). Next, we describe these methods in more detail, and Fig. 5 illustrates some of these methods.
  1. (1)
    Translation: This method adds a random constant displacement value to both x and y coordinates while ensuring the distance between coordinates remains the same (Wang et al., 2022). All points will be shifted to a fixed distance and direction.
  2. (2)
    Flipping: This method manipulates the orientation of axes, both on a local and global scale. There are several flipping techniques, such as global horizontal flipping, which flips the map’s horizontal axis; global vertical flipping, which flips the map’s vertical axis; global axes flipping, which flips both horizontal and vertical axes; and local random flipping, which randomly alters the orientation of horizontal, vertical, or both axes within grid cells to hide location points (Gupta and Rao, 2020; Leitner and Curtis, 2004, 2006).
  3. (3)
    Rotating: This method rotates each point by a fixed angle ($0\leq\theta\leq 360$), centering the rotation around either the origin of the coordinate system or an arbitrarily chosen point (Wang et al., 2022).
  4. (4)
    Scaling (Resizing): This method multiplies a constant value with both x and y coordinates, which remains consistent for both axes (Wang et al., 2022). Consequently, this process alters the distances between the points and all points will be expanded by a scaling factor.
  5. (5)
    Horizontal shear: This method employs a linear transformation to horizontally shear the location points (Armstrong et al., 1999; Broen et al., 2021). Each point is shifted along its x-axis, oriented at a 45° angle relative to its initial position concerning the center of the point distribution.
Affine transformations preserve the spatial structure of the data, proving useful for spatial analysis and visualisation methods such as clustering, while maintaining the number of records and avoiding random relocation of data (Ajayakumar et al., 2019). However, in these techniques, if the transformation matrix is known, retrieving the coordinates of the original records from the perturbed dataset is simple. Therefore, affine transformations are not a widely used methods (Wang et al., 2022).
@@ 3.1.4. Spatial aggregation & Clustering
Spatial aggregation and clustering techniques are commonly used to protect individual privacy by reducing spatial resolution, making it difficult to pinpoint exact locations and preventing the disclosure of precise geographical information. However, these techniques can also result in a loss of data granularity, which may restrict detailed analysis and reduce overall location accuracy.
Spatial aggregation aggregates individual location points within specific geographic units or areas, such as cities, zipcodes, or administrative areas, to ensure anonymity and prevent the disclosure of exact locations. This technique operates by decreasing the spatial resolution of the data, thereby modifying the original data patterns. The goal is to generate generalised records that consolidate information within defined territories, protecting the identities of specific locations. However, this process may compromise the utility of published records as the detailed patterns found in high-resolution data might not be as discernible in the aggregated information. Figure 6 illustrates different spatial aggregation techniques.
Clustering aims to reveal inherent patterns within spatial data by grouping similar data points based on geographical proximity or shared characteristics. Clustering maintains the overall characteristics of an area without pinpointing precise locations, preserving the general patterns while ensuring individual data privacy. However, due to the computations applied directly on original data, clustering could expose sensitive information especially in datasets containing personal or confidential data. This can lead to unintended re-identification of individuals within the dataset, posing risks of privacy breaches and discrimination. Adversaries can exploit clustering results to deduce individuals’ group memberships or glean details about the data distribution, even without direct access to the data (Canbay and Sever, 2015).
@@ 3.1.4.1 Grid-based aggregations
The techniques in this category involve the realignment of individual data points onto a grid system (Leitner and Curtis, 2004). It typically involves relocating each original data point to specific positions within designated grid cells. These relocations may take place at either the centroid or vertices of the grid cells, each being aligned with a defined side length. The efficacy of methods within this category is notably influenced by two crucial parameters: (1) the size of the aggregation units and (2) the boundaries utilised to define these units.
In grid-based aggregations, the grid cell size could be fixed or varied. Grid line masking and grid center masking are two popular methods that use fixed-size grid cells. The grid line masking method relocates location points to the nearest edge of the grid cell in which they are contained (Gupta and Rao, 2020). The grid center masking method relocates the location points to the centroid of the cell in which they are presently situated (Leitner and Curtis, 2004).
@@
    Quadtree hierarchical geographic data structures is a method that uses varying-size grid cells. This method employs a quadtree methodology on an initial 1km2 European grid and population data points to create a grid with varying sizes that can be adjusted according to local population density(Lagonigro et al., 2017). In this approach, densely populated areas are divided into smaller square areas to ensure accurate data dissemination, while sparsely populated areas are grouped into larger areas to prevent the identification of individual data.
@@
    Military Grid Reference System (MGRS) protects sensitive point-based geospatial data with a focus on preventing the matching of points with their true underlying observations, such as street addresses. This method involves converting coordinates to the Military Grid Reference System (MGRS) and using digit switching for masking, allowing encryption at different spatial precision levels (Clarke, 2016). The selection of the digit switching combination minimises differences in aggregate descriptive spatial statistics between the masked and unmasked data, with emphasis on the nearest neighbour statistic. Despite the large runtime, this method protect sensitive geospatial data, overcoming the disadvantages of geometric point masking methods, such as the lack of control over parameters, compromised spatial analysis of distribution, and susceptibility to violations of geoprivacy associated with the use of high-resolution imagery and accurate data from Global Navigation Satellite Systems (GNSS). The proposed digit-switching method, which operates effectively across multiple nested spatial scales, addresses these limitations by allowing the users to choose a distribution that closely mimics the spatial distribution of unmasked data, ensuring suitability for spatial analysis based on aggregate descriptive statistics.